### v1

# A look at the operator product expansion in critical dynamics

Abstract
We consider the critical relaxation of the Ising model, the so-called model A, and study its operator product expansion. Within perturbation theory, we focus on the operator product expansions of the two-point function and the response function. At the fixed point, we normalize the coefficients and the scaling variables so that the result displays universality. The role of the fluctuation-dissipation theorem is also discussed, and it is shown that it provides non-perturbative relations among the operator product expansion coefficients. Finally, the large $N$ limit is considered.
I Introduction
The operator product expansion (OPE) offers a non-perturbative tool to study quantum and statistical field theories. The paradigmatic example is given by two-dimensional critical phenomena, which have been studied in detail also thanks to the application of conformal field theory (CFT) techniques in which the OPE plays a key role Francesco et al. (1997). More recently, the conformal bootstrap approach allowed to obtain precise predictions also in higher dimensions Poland et al. (2019). These results have been obtained by exploiting also the powerful constraints due to conformal symmetry, which, however, may not always be present since, for instance, one may not exactly be at criticality or because of the presence of some external scale, as it happens in certain non-equilibrium systems.
However, as such, the OPE holds also outside the critical region, i.e., away from a fixed point, and it can be employed to study a wide variety of systems. In particular, the OPE is used to study the expansion of the two-point function of equilibrium systems close to criticality, see, e.g., Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015). In the context of QCD, it is employed to study deep inelastic scattering processes Novikov et al. (1978).
The purpose of this paper is to study the operator product expansion in the context of critical dynamics in the case where the fluctuation-dissipation theorem applies for the first time. We shall show that the OPE allows one to define further universal quantities, besides the critical exponents, which generalize the OPE coefficients of equilibrium statistical field theories at criticality. We shall work within the response field formalism Martin et al. (1973); Janssen (1976); de Dominicis (1976), in which the so-called response field $\bar{\phi}$ is introduced over and above the physical field $\phi$. This makes the formalism very similar to what is employed in genuinely non-equilibrium systems and so it constitutes a first step towards a general study of the OPE in non-equilibrium systems. In this respect, it should be noted that a sort of short-time expansion is sometimes invoked in the studies concerning the so-called critical initial slip, which is related to initial states that are not the equilibrium ones, see Täuber (2014) and references therein. Furthermore, as time and space scale differently, our results are somehow reminiscent of the OPE for non-relativistic theories, see Golkar and Son (2014); Shimada and Shimada (2021).
The paper is organized as follows. In section II we introduce our notation and conventions and briefly review some of the main features of the path integral approach to model A. In section III we introduce some composite operators of interest and analyze their renormalization and expectation value (VEV). We also briefly consider the consequences of the time-reversal symmetry at the origin of the fluctuation-dissipation theorem. In section IV we derive the Gaussian, i.e., free-field theory, OPE. We do so by considering the functional “Schroedinger” picture associated with the Fokker-Planck equation. In section V we compute the leading order corrections to the most singular parts of the OPE of the two-point and the response functions. Finally, in section VI we discuss how to extract universal information from the OPE calculated in the previous section and derive a universal form of the OPE coefficients to $O\left(\epsilon\right)$. In sections VII and VIII we consider the $O\left(N\right)$ model both perturbatively and in the large $N$ limit. We summarize our findings in section IX. Some definitions and technical discussions are confined to Appendices A and B.
II Path integral approach to model A
In this section, we state our conventions and recall the main features of model A. For a review, we refer to Zinn-Justin (2002).
II.1 Model A and its path integral representation
The dynamics of model A is defined by the following stochastic differential equation (SDE)
$\partial_{t}\phi_{0}\left(t,x\right)$ $=$ $-D_{0}\frac{\delta H}{\delta\phi_{0}\left(t,x\right)}+\zeta\left(t,x\right)\,,$ (1)
with $\zeta$ being a Gaussian noise with zero average and covariance given by
$\langle\zeta\left(t,x\right)\zeta\left(t^{\prime},x^{\prime}\right)\rangle$ $=$ $2D_{0}\delta\left(t-t^{\prime}\right)\delta\left(x-x^{\prime}\right)\,.$
We consider the SDE (1) in the Ito sense. The Hamiltonian $H$ is defined to be that of the Landau-Ginzburg model:
$H$ $=$ $\int_{x}\left[\frac{1}{2}\partial_{i}\phi_{0}\partial_{i}\phi_{0}+\frac{1}{2}m_{0}^{2}\phi_{0}^{2}+\frac{1}{4!}g_{0}\phi_{0}^{4}\right]\,,$ (2)
where $\partial_{i}\equiv\partial/\partial x^{i}$. The zero in the subscript of both couplings and fields signals that those are bare (non-renormalized) quantities. Our conventions and notations for the integrals and the Fourier transform are given in Appendix A.
Associated with the stochastic process that describes the dynamics of model A, i.e., the SDE (1), there is the following Fokker-Planck equation Zinn-Justin (2002):
$\partial_{t}P\left[\phi\right]\left(t\right)$ $=$ $D\Biggr{\{}\int_{x}\frac{\delta}{\delta\phi\left(x\right)}\left(\frac{\delta H}{\delta\phi\left(x\right)}\right)+\int_{x}\frac{\delta^{2}}{\delta\phi\left(x\right)^{2}}\Biggr{\}}P\left[\phi\right]\left(t\right)$ (3)
$\equiv$ $-H_{{\rm FP}}P\left[\phi\right]\left(t\right)\,,$
which implicitly defines the “Fokker-Planck Hamiltonian” $H_{{\rm FP}}$.
The stochastic process (1) has also an associated path integral representation, which has been introduced thanks to the works of Martin, Siggia, Rose, Janssen, and De Dominicis Martin et al. (1973); Janssen (1976); de Dominicis (1976). Such a path integral is constructed by introducing a further field, which is referred to as the “response field” and is denoted by $\bar{\phi}$. Thus, one introduces the following generating functional (see, e.g., Täuber (2014)):
$Z\left[J_{\phi},J_{\bar{\phi}}\right]$$\equiv$ $\int{\cal D}\phi{\cal D}\bar{\phi}\,\exp\left[-S\left[\phi,\bar{\phi}\right]+\int_{t,x}J_{\phi}\left(t,x\right)\phi\left(t,x\right)+\int_{t,x}J_{\bar{\phi}}\left(t,x\right)\bar{\phi}\left(t,x\right)\right]\,,$
where the action is given by
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}\left(t,x\right)\left(\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\frac{\delta H}{\delta\phi_{0}\left(t,x\right)}-D_{0}\bar{\phi}_{0}\left(t,x\right)^{2}\right)\,.$ (4)
It is important to note that the above action is invariant under the following time reversal symmetry (TRS) transformation:
$\begin{cases}\phi_{0}\left(t,x\right)\rightarrow\phi_{0}\left(-t,x\right)\\ \bar{\phi}_{0}\left(t,x\right)\rightarrow\bar{\phi}_{0}\left(-t,x\right)-\frac{1}{D_{0}}\dot{\phi}_{0}\left(-t,x\right)\end{cases},$ (5)
where the dot indicates a time derivative. This symmetry implies the so-called fluctuation-dissipation theorem (FDT), which relates the two-point function with the response function. For a more detailed discussion of these aspects, we refer the reader to Täuber (2014); Canet and Chate (2007); Sieberer et al. (2015).
II.2 Perturbative renormalization of model A in $4-\epsilon$ dimensions
By inserting the Hamiltonian (2) in (4) one obtains the following action
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}\left(t,x\right)\Biggr{\{}\partial_{t}\phi_{0}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}\left(t,x\right)^{2}\right)\phi_{0}\left(t,x\right)$ (6)
$-D_{0}\bar{\phi}_{0}\left(t,x\right)^{2}\Biggr{\}}\,.$
By expressing (6) in terms of the renormalized parameters one writes
$S$ $=$ $\int_{t,x}Z_{\bar{\phi}}^{1/2}Z_{\phi}^{1/2}\bar{\phi}\left(t,x\right)\Biggr{\{}\partial_{t}\phi\left(t,x\right)+Z_{D}D\left(-\partial^{2}+Z_{m^{2}}m^{2}+\frac{Z_{g}g}{3!}Z_{\phi}^{1/2}\phi\left(t,x\right)^{2}\right)\phi_{0}\left(t,x\right)$ (7)
$-Z_{\bar{\phi}}Z_{D}D\bar{\phi}_{0}\left(t,x\right)^{2}\Biggr{\}}\,.$
The model is perturbatively renormalizable Täuber (2014) and we shall work with the minimal subtraction scheme in dimensional regularization so that the counterterms $Z_{i}-1$ can be expressed as a sum of poles in $\epsilon$. It should be noted that the renormalization factors $Z_{i}$ are not fully independent from one another, they are related by the fluctuation-dissipation theorem, i.e., the symmetry (5). The renormalization factors appearing in the action (7) have been calculated many times and we refer the reader to Täuber (2014) for a review.
Moreover, since in the static limit one recovers the equilibrium distribution, the renormalization, and so the critical exponents, of the couplings appearing in the Hamiltonian (2) are the same as in the static case. However, when discussing dynamical critical systems, a further exponent arises, the so-called dynamical critical exponent $z$, which describes the relation between spatial and temporal scalings and is related to the renormalization of the coupling $D$: $z=2+\gamma_{D}$ with $\gamma_{D}=D^{-1}\mu\frac{d}{d\mu}{D}=-Z_{D}^{-1}\mu\frac{d}{d\mu}{Z}_{D}$. At the Gaussian fixed point one has $z=2$, which implies that a two-point function takes the following scaling form: $G\left(t,x\right)=x^{-2\Delta}g\left(\frac{t}{x^{2}}\right)$.
Finally, let us assign a (bare) scaling dimension to each field. The scaling dimension of the field $\phi$ is $\frac{d-2}{2}$, i.e., $\phi$ scales like $x^{-\frac{d-2}{2}}$ while the scaling dimension of the response field $\bar{\phi}$ is $\frac{d+2}{2}$, i.e., $\bar{\phi}$ scales like $x^{-\frac{d+2}{2}}$.
III Composite operators
In this section we consider the first order renormalization of some composite operators, which we shall use later on to construct the OPE. We also determine the normalizations of interests.
Let us start by making few general considerations. We shall work in $d=4-\epsilon$, within perturbation theory, and adopt dimensional regularization together with minimal subtraction to renormalize composite operators. In this setting, in order to define a composite operator $\left[O\right]$, one has that the “naive” operator $O$ requires mixing with operators of the same (or smaller) scaling dimension, schematically $\left[O_{i}\right]=Z_{ij}O_{j}$. For example, the operator $\phi^{2}$ can mix only with the identity, $m^{2}\mathbb{I}$, whereas the operator $\phi^{4}$ mixes with $\phi\bar{\phi}$, among other operators. This latter example tells us that, in the dynamical extension of the Landau-Ginzburg Hamiltonian, the composite operators require new mixing with respect to the static case, which clearly does not possess any mixing with operators depending on the response field.111 Let us note that in principle one could integrate out the response field and work with an Onsager-Machlup kind of action Onsager and Machlup (1953). In this setting, one would not have any mixing with $\bar{\phi}$-dependent operators since the response field has been integrated out. However, throughout this work, we shall always keep the response field since we are interested also in the response function.
III.1 The composite operator $\left[\phi^{2}/2\right]$
We wish to construct the renormalized composite operator $\left[\phi^{2}/2\right]$ to order $O\left(\epsilon\right)$. The composite operator takes the form:
$\left[\frac{\phi^{2}}{2}\right]$$\equiv$ $Z_{22}Z_{\phi}\frac{\phi^{2}}{2}+Z_{20}m^{2}\,,$
where actually the wave function renormalization is negligible at this order.
At the Gaussian level, only the VEV of the operator requires renormalization. By demanding that $\langle\left[\phi^{2}/2\right]\rangle$ be finite one obtains
$Z_{20}$ $=$ $\frac{1}{16\pi^{2}\epsilon}\,.$
The diagram associated with $\langle\left[\phi^{2}/2\right]\rangle$ is shown in Figure 1.
Figure 1: The diagram 0(a) is the diagram contributing to the VEV of $\left[\phi^{2}/2\right]$ at Gaussian order. The diagrams 0(b) contributes at order $O(g)$ while diagram 0(c) vanishes. We do not display here the diagrams associated with the counterterms stemming out of the renormalization factors $Z_{22}$, $Z_{20}$, and $Z_{m^{2}}$.
In order to have an idea about what kind of expression the VEV of $\left[\phi^{2}/2\right]$ has, let us write it down explicitly within the context of dimensional regularization:
$\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}\,=\,\Bigr{\langle}\left(\frac{\phi^{2}}{2}+\frac{m^{2}\mu^{-\epsilon}}{16\pi^{2}\epsilon}\right)\Bigr{\rangle}+O\left(g\right)$ $=$ $\frac{m^{2}}{32\pi^{2}}\left(\gamma-1+\log\frac{m^{2}}{4\pi\mu^{2}}\right)+O\left(\epsilon\right)\,,$
where we do not display explicitly the $O\left(\epsilon\right)$-terms, which however should be kept.
The mixing matrix $Z_{ij}$ can be expanded in a power series with respect to the coupling constant $g$. We employ the following notation:
$Z_{ij}$ $=$ $\sum_{k}Z_{ij,k}g^{k}\,.$
At the Gaussian level, one finds $Z_{22,0}=1$ and $Z_{20,0}=\frac{1}{16\pi^{2}\epsilon}$.
To order $O\left(g\right)$ one obtains the following result for the counterterms
$Z_{22,1}$ $=$ $\frac{1}{\left(4\pi\right)^{2}}\frac{1}{\epsilon}$
$Z_{20,1}$ $=$ $\frac{1}{256\pi^{4}\epsilon^{2}}\,.$
The counterterm $Z_{22,1}$ is determined by considering the diagram in Figure 2.
Figure 2: The diagram is associated with the three-point function $\langle\phi\phi\left[\frac{\phi^{2}}{2}\right]\rangle$ and determines $Z_{22,1}$.
The counterterm $Z_{22,1}$ implies the anomalous dimension $\gamma_{\phi^{2}}=-Z_{\phi^{2}}^{-1}\dot{Z}_{\phi^{2}}\approx\frac{g}{16\pi^{2}}$. At the fixed point, one obtains the following correction to the scaling dimension of the operator:
[…]
where we used the fixed point value $g_{*}\approx 16\pi^{2}\epsilon/3$.
The $O\left(g\right)$-correction to the VEV $\langle\left[\phi^{2}/2\right]\rangle$ is determined by the diagrams shown in Figure 1 together with those involving the relevant counterterms. In the presence of translational invariance, which applies to our setting, we expect the VEV to be the same as in the static case and not to depend on the coupling $D$. It can be checked indeed that the VEV is the same in the static and in the dynamical theory.
Finally, let us also consider the static limit of the two-point of $\left[\phi^{2}/2\right]$ to $O\left(g\right)$. At the fixed point, the diagrams contributing to this order are shown in Figure 3. For our purposes, we shall be interested in defining a normalization constant ${\cal N}_{2}$ such that the operator $\phi_{2}\equiv{\cal N}_{2}\left[\frac{\phi^{2}}{2}\right]$ is normalized as follows
$\Bigr{\langle}\phi_{2}\left(0,x\right)\phi_{2}\left(0,0\right)\Bigr{\rangle}\,=\,{\cal N}_{2}^{2}\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\left(0,x\right)\left[\frac{\phi^{2}}{2}\right]\left(0,0\right)\Bigr{\rangle}$ $=$ $\frac{1}{x^{2\Delta_{2}}}\,.$
Figure 3: The diagrams contributing to $\langle\left[\frac{\phi^{2}}{2}\right]\left(t,x\right)\phi\left[\frac{\phi^{2}}{2}\left(0,0\right)\right]\rangle$ up to $O(g)$. The diagrams associated with the counterterms are not displayed.
After an explicit computation and performing the $\epsilon$-expansion at the fixed point, one obtains:
${\cal N}_{2}$$\approx$ $4\sqrt{2}\pi^{2}-\epsilon\frac{2}{3}\sqrt{2}\left(2\gamma\pi^{2}-\pi^{2}+2\pi^{2}\log\pi-2\pi^{2}\log\mu\right)\,.$
III.2 A consequence of the time-reversal symmetry
Let us show that the renomalized composite operator $\left[\bar{\phi}\phi\right]$ is related to $\left[\phi^{2}/2\right]$ by the time reversal symmetry (5). As a first step, let us consider the transformation of the monomial $\phi_{0}\bar{\phi}_{0}$:
$\phi_{0}^{\prime}\left(t,x\right)\bar{\phi}_{0}^{\prime}\left(t,x\right)$ $=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)-\frac{1}{D_{0}}\phi_{0}\left(-t,x\right)\dot{\phi}_{0}\left(-t,x\right)$
$=$ $\phi_{0}\left(-t,x\right)\bar{\phi}_{0}\left(-t,x\right)+\frac{1}{D_{0}}\partial_{t}\left(\frac{\phi_{0}\left(-t,x\right)^{2}}{2}\right)\,.$
The last term on the RHS takes almost the form of the composite operator $\partial_{t}\left[\phi^{2}/2\right]$. If we multiply the previous expression by the suitable renormalization factor, i.e., $Z_{22}$, one finds
$Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi^{\prime}\left(t,x\right)\bar{\phi}^{\prime}\left(t,x\right)$ $=$ $Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)+\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\,.$
Thus, if we consider the TRS in a correlation function one obtains
$\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)\Psi\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}$ $=$ $\Bigr{\langle}Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(-t,x\right)\bar{\phi}\left(-t,x\right)\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}$ (8)
$+\Bigr{\langle}\frac{1}{D}\partial_{t}\left[\frac{\phi\left(-t,x\right)^{2}}{2}\right]\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)\Bigr{\rangle}\,,$
where $\Psi\left(\left\{t_{i},x_{i}\right\}\right)$ denotes a bunch of $\phi$- and $\bar{\phi}$-fields and $\Psi_{{\rm TRS}}\left(\left\{t_{i},x_{i}\right\}\right)$ its transformation under TRS. The last correlation in (8) is finite since it involves renormalized fields and a renormalized composite operator. The other two correlation functions involve the same operator, $Z_{D}Z_{22}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\bar{\phi}$, evaluated at different times and inserted in different correlation functions. By assuming the absence of non-trivial systematic cancellations between these two correlation functions it follows that also the first two correlations appearing in (8) are each finite on their own. In turn, this implies that the renormalization factor $Z_{D}Z_{22}$ is the additional renormalization required by $\left[\phi\bar{\phi}\right]$:
$\left[\phi\bar{\phi}\right]$ $=$ $Z_{\phi\bar{\phi}}Z_{\phi}^{1/2}Z_{\bar{\phi}}^{1/2}\phi\left(t,x\right)\bar{\phi}\left(t,x\right)$
with $\mbox{with }Z_{\phi\bar{\phi}}=Z_{D}Z_{22}\,.$
Actually, the assumption of absence of cancellations in (8) is not necessary. Indeed, it is possible to check that $\left[\phi\bar{\phi}\right]$ cannot mix with any other operator.
At the fixed point, the TRS entails a relation between the scaling dimension of $\left[\phi^{2}/2\right]$ and $\left[\phi\bar{\phi}\right]$. Given $Z_{\phi\bar{\phi}}=Z_{D}Z_{22}$, it is straightforward to check that
$\gamma_{\phi^{2}}\,=\,-Z_{22}^{-1}\dot{Z}_{22}$ $=$ $2-z+\gamma_{\phi\bar{\phi},\phi\bar{\phi}}\,,$
where $\gamma_{\phi\bar{\phi},\phi\bar{\phi}}\equiv-Z_{\phi\bar{\phi}}^{-1}\dot{Z}_{\phi\bar{\phi}}$. Hence the following relation between the scaling dimensions holds:
$\Delta_{\phi\bar{\phi}}$ $=$ $d_{\phi\bar{\phi}}+\gamma_{\phi\bar{\phi},\phi\bar{\phi}}$ (9)
$=$ $d_{\phi^{2}}+\gamma_{\phi^{2}}+z\,=\,\Delta_{\phi^{2}/2}+z\,.$
At order $O\left(g\right)$ it is straightforward to check that $\gamma_{\phi\bar{\phi},\phi\bar{\phi}}=\gamma_{\phi^{2}}+O\left(g^{2}\right)$ with $\gamma_{D}=O\left(g^{2}\right)$. The relation (9) has also been discussed in Calabrese and Gambassi (2004), to which we refer for a more detailed discussion regarding the TRS.
IV Gaussian operator product expansion
In this section, we consider the OPE for the Gaussian theory. As the theory is Gaussian one can check the expected results in different ways. We found it instructive to calculate the OPE expansion by developing a sort of normal ordering prescription, thus mimicking standard quantum field theory in which the Gaussian OPE can be derived from the Wick’s theorem Bonneau (2009). In particular, we shall see that the following statements hold:
$\phi\left(t,x\right)\phi\left(0,0\right)$ $=$ $G_{\phi\phi}\left(t,x\right)+\left[\phi^{2}\right]+\mbox{higher derivatives}$ (10)
$\phi\left(t,x\right)\bar{\phi}\left(0,0\right)$ $=$ $G_{\phi\bar{\phi}}\left(t,x\right)+\left[\phi\bar{\phi}\right]+\mbox{higher derivatives}\,,$ (11)
where $\left[\phi^{2}\right]$ and $\left[\phi\bar{\phi}\right]$ are defined in terms of the normal ordering prescription, which we shall introduce later on. Thus, the RHS of equations (10) and (11) define the OPE coefficients associated with $\phi\times\phi$ and $\phi\times\bar{\phi}$. For instance, one recognizes that the most singular coefficient of the $\phi\times\phi$ OPE is associated with the identity operator, which we shall denote by $c_{\phi\phi 1}$ with $c_{\phi\phi 1}=G_{\phi\phi}\left(t,x\right)$ in the present case. More generally, we shall denote the coefficient associated with an operator $O$ in the product $O_{1}\times O_{2}$ by $c_{O_{1}O_{2}O}$. Finally, let us note the absence in the $\phi\times\bar{\phi}$ OPE of the operator $\left[\phi^{2}\right]$ at the Gaussian level.
IV.1 Gaussian Fokker-Planck Hamiltonian
We regard the Fokker-Planck equation (3) as a sort of (functional) Schroedinger equation with an imaginary time. This allows us to work in the Schroedinger picture and eventually move to the Heisenberg picture. We refer the reader to Hatfield (1992) for an overview of the Schroedinger picture of QFT and to Sakita (1983) for its application to the Fokker-Planck equation in connection to the stochastic quantization program. As a first step, we introduce the following analog of the position and momentum operators:
$\hat{\phi}\left(x\right)$ $=$ $\phi\left(x\right)$
$\hat{\pi}\left(x\right)$ $=$ $\frac{1}{i}\frac{\delta}{\delta\phi\left(x\right)}\,.$
By using the above operators, the Fokker-Planck equation for the free theory can be written as follows:
$\partial_{t}P\left[\phi\right]\left(t\right)$ $=$ $D\Biggr{\{}\int_{x}i\hat{\pi}\left(x\right)\left(-\partial_{x}^{2}+m^{2}\right)\hat{\phi}\left(x\right)-\int_{x}\hat{\pi}\left(x\right)^{2}\Biggr{\}}P\left[\phi\right]\left(t\right)\,.$
Note that the first term is non-Hermitian. Therefore, the “Hamiltonian” on the RHS is not Hermitian as well. [^2] Nevertheless, it is interesting to note that $H_{{\rm FP}}$ appears to be “PT-symmetric” in the sense that it is invariant under $\pi\rightarrow\pi$, $\phi\rightarrow-\phi$, and $i\rightarrow-i$. Next, we introduce the following operators:
$\hat{a}\left(p\right)$$\equiv$ $\beta\Delta_{p}\left(\hat{\phi}\left(p\right)+i\frac{1}{\Delta_{p}}\hat{\pi}\left(p\right)\right)$ (12)
$\hat{b}\left(p\right)$$\equiv$ $\frac{1}{i\beta\Delta_{p}}\hat{\pi}\left(-p\right)\,,$ (13)
where $\beta$ is an arbitrary parameter and $\Delta_{p}\equiv\left(p^{2}+m^{2}\right)$. The operators $\hat{a}$ and $\hat{b}$ satisfy the following algebra:
$\left[\hat{a}\left(p\right),\hat{b}\left(q\right)\right]$ $=$ $\left(2\pi\right)^{d}\delta\left(p-q\right)\,,$
which is the same as that of creation-annihilation operators (see, e.g., Peskin and Schroeder (1995)) but with $b$ being replaced by $a^{\dagger}$.
A straightforward calculation shows then that the Fokker-Planck Hamiltonian takes the following form
$H_{{\rm FP}}$ $=$ $D\int_{k}\Delta_{k}\left(\hat{b}\left(k\right)\hat{a}\left(k\right)\right)\,=\,\int_{k}\omega_{k}\left(\hat{b}\left(k\right)\hat{a}\left(k\right)\right)\,,$ (14)
with $\omega_{k}\equiv D\Delta_{k}$. Again, let us note that since both $\hat{a}$ and $\hat{b}$ are not Hermitian also $H_{{\rm FP}}$ is not.
IV.2 Normal ordering
One can re-express the original fields $\phi$ and $\bar{\phi}$ in terms of the creation and annihilation operators (1213). The Heisenberg operators are defined by
$\phi\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\phi\left(x\right)e^{-H_{{\rm FP}}t}$
$\bar{\phi}\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\bar{\phi}\left(x\right)e^{-H_{{\rm FP}}t}\,.$
Finally, let us introduce the normal ordering operation, which we denote by sandwiching any operator in colons, i.e., $:\hat{O}:$. We define normal ordering as the operation that maps a given operator $\hat{O}\left(\hat{a},\hat{b}\right)$ to a new operator obtained from $\hat{O}$ by moving the creation and annihilation operators so that the $\hat{b}$-operators are on left of the $\hat{a}$-operators. For example, the Hamiltonian (14) is normal ordered.
It is then possible to follow a textbook presentation of the Wick’s theorem and check that the following statements holds:
$T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)$ $=$ $G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{1},x_{1}\right)\phi\left(t_{2},x_{2}\right):$
$T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right):\,,$
where the $T$ on the left hand side denotes the time ordering.
At this point, the OPE amounts to the Taylor expansion of the normal ordering statement:
$T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)$ $=$ $G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)^{2}:+\cdots$
$T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)\bar{\phi}\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right):+\cdots\,.$
It should be noted that, strictly speaking, the normal ordering represents one possible regularization for the composite operators of the free theory. In particular, the expectation value on any normal ordered operator vanishes by construction. However, this is just a possible choice and other choices, such as dim-reg discussed in section III.1, lead to different values of the VEV. Clearly, also the associated OPE coefficients are scheme dependent in general. However, one can obtain universal and scheme independent results, as we shall check explicitly at the fixed point.
Finally, let us consider the OPE of $\bar{\phi}\times\bar{\phi}$, at the Gaussian level one has
$T\bar{\phi}\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $:\bar{\phi}\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)^{2}:+\cdots\,.$
It is also straightforward to check that $c_{\bar{\phi}\bar{\phi}1}=c_{\bar{\phi}\bar{\phi}\left[\phi^{2}\right]}=0$ exactly. To see this, first consider the OPE within $0=\langle\bar{\phi}\bar{\phi}\rangle$: this relation implies that only operators having vanishing VEVs appear in the OPE. In particular $c_{\bar{\phi}\bar{\phi}1}=0$. Next, let us consider the OPE of $\bar{\phi}\times\bar{\phi}$ in $0=\langle\bar{\phi}\bar{\phi}\bar{\phi}\bar{\phi}\rangle$. Given $c_{\bar{\phi}\bar{\phi}1}=0$, the most singular contribution comes from $0=\langle c_{\bar{\phi}\bar{\phi}\left[\phi^{2}\right]}\left[\phi^{2}\right]\bar{\phi}\bar{\phi}\rangle$. Since $\langle\left[\phi^{2}\right]\bar{\phi}\bar{\phi}\rangle\neq 0$ at tree level one must have $c_{\bar{\phi}\bar{\phi}\left[\phi^{2}\right]}=0$.
V Leading corrections to the OPE
In this section we consider the OPE associated with the products $\phi\times\phi$ and $\phi\times\bar{\phi}$ and compute the most singular contributions to the first order in perturbation theory. We shall focus on deriving the OPE coefficients at the fixed point. Let us recall that the OPE applies to the short-distance limit. In our case, for instance, we define the short-distance limit of $\phi\left(t,x\right)\phi\left(0,0\right)$ as the limit $x\rightarrow 0$ with $t/x^{z}$ being held fixed.
One may adopt various strategies to compute OPE coefficients: for instance, one may look at the three- or four-point function and consider the short distance limit of two of the operators appearing in such correlations, see, e.g., Collins (2023). In this section we follow a different approach, based on the mass expansion of the two-point function, which leads to the same result, as we have checked explicitly. On the one hand, this approach allows one to derive the OPE coefficients and at the same time to calculate the two-point function close to criticality, i.e., for values of $m$ small but non-zero. Indeed, one of the first applications of the OPE was precisely the prediction of the form of the static two-point function close to criticality in the regime $m\ll p$ Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015).
As we already mentioned, we shall work in the MS scheme. This implies that the OPE coefficients are analytic in the mass, see Tkachov (1983); Guida and Magnoli (1996) and references therein. It follows that the non-analytic dependence of the two-point function on the mass is entirely due to the VEV of the operators once the OPE has been performed. A key point here is also that the most singular OPE coefficients are multiplied by the operators that have the smallest mass dimension. For instance, in 4D one expects $\langle\phi^{2}\rangle\sim m^{2}$ while $\langle\phi^{4}\rangle\sim m^{4}$. Of course, fluctuations modify the exponents as dictated by the renormalization group. Let us consider the product $\phi\times\phi$ as an example, one has
$\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$$\approx$ $\left(c_{\phi\phi 1}^{\left(0\right)}\left(t,x\right)+m^{2}c_{\phi\phi 1}^{\left(1\right)}\left(t,x\right)+\cdots\right)+\left(c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)+\cdots\right)\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}$ (15)
$+o\left(m^{2}\right)\,,$
where $\langle\left[\phi^{2}\right]\rangle\sim m^{2}$ and the superscript $(n)$ in $c^{(n)}$ denotes the order of the expansion in the mass paramter. By expanding the LHS of (15) one can calculate the OPE coefficients since they are the only unknowns after inserting the value calculated for the VEVs. We remark that the coefficients $c_{\phi\phi 1}^{\left(0\right)}\left(t,x\right)$ and $c_{\phi\phi\left[\phi^{2}\right]}^{\left(0\right)}\left(t,x\right)$ evaluated at $g=g_{*}$ are actually the fixed point coeffients, which is our main interest.
V.1 OPE expansion of $\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$
In section III we have computed the VEV of the operator $\left[\phi^{2}/2\right]$ up to order $O\left(g\right)$. In order to apply the strategy outlined at the beginning of this section, we must compute the LHS of equation (15) up to the same order. At the tree level the two-point function is given by equation (23). The diagrams contributing to order $O\left(g\right)$ only involve the mass renormalization. Thus, it turns out that the LHS of (15) is given precisely by (23) once the substitution
$m^{2}$$\rightarrow$ $m^{2}+\frac{1}{2}g\frac{m^{2}}{\left(4\pi\right)^{2}}\left(\log\left(\frac{m^{2}}{4\pi\mu^{2}}\right)+\gamma-1\right)$ (16)
is made and the expansion to first order in $g$ is performed. (We work to the leading order in $\epsilon$.)
Let us note that in the present case $\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$ is not at the fixed point: we shall set $g=g_{*}$ but keep a non-zero, albeit small, mass so that we are a bit off criticality. By matching each term in the expansions of the LHS and RHS of equation (15), one obtains the following results:
$c_{\phi\phi 1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\frac{1}{x^{d-2}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}+O\left(\epsilon^{2}\right)$
$c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $2+g_{*}\frac{1}{16\pi^{2}}\Biggr{\{}\log\left(\pi x^{2}\mu\right)+\gamma+\Gamma\left(0,\frac{x^{2}}{4Dt}\right)\Biggr{\}}+O\left(\epsilon^{2}\right)\,,$
[…]
Let us remark that the same result for the OPE coefficients can be obtained by looking at the four-point function, similarly to what is done in Collins (2023).
V.2 OPE expansion of $\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$
We consider the OPE expansion of $\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$, i.e, the response function. In analogy with the reasoning of section V.1, we first consider the mass expansion of the response function to order $O\left(g\right)$. To do so, we consider the Gaussian expression, equation (24). The order $O\left(g\right)$ is reproduced by implementing the substitution (16) and expanding in the coupling constant. Then we consider
$\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$$\approx$ $\left(c_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)+m^{2}c_{\phi\bar{\phi}1}^{\left(1\right)}\left(t,x\right)+\cdots\right)+\left(c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)+\cdots\right)\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}$
$+o\left(m^{2}\right)\,.$
The LHS is proportional to $\theta\left(t\right)$ and so are the OPE coefficients on the RHS.
By matching the terms on the LHS and RHS, we deduce
$c_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{\theta\left(t\right)}{\left(4\pi\right)^{d/2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)^{d/2}}+O\left(\epsilon^{2}\right)$
$c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $-g_{*}\frac{\theta\left(t\right)}{\left(4\pi\right)^{2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)}+O\left(\epsilon^{2}\right)\,,$
where we did not expand $c_{\phi\bar{\phi}1}^{\left(0\right)}$ in $\epsilon$ while the coefficient $c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}$ is already expressed to $O\left(\epsilon\right)$.
V.3 Relation to the fluctuation-dissipation theorem
The time-reversal symmetry (4) implies exact relations among various correlation functions. It would be interesting to study in detail the consequence of the symmetry (4) for the OPE. In this section we limit ourselves to the following consideration. Let us recall the fluctuation-dissipation theorem:
$\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$ $=$ $-\frac{1}{D}\partial_{t}\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle\,,$ (17)
where $t>0$ and we work in the MS-scheme. We can expand via the OPE each of the correlation functions appearing in (17). One obtains
$\Bigr{\langle}c_{\phi\bar{\phi}1}\left(t,x\right)+c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)\left[\frac{\phi^{2}}{2}\right]+\cdots\Bigr{\rangle}$ $=$ $-\frac{1}{D}\partial_{t}\Bigr{\langle}c_{\phi\phi 1}\left(t,x\right)+c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)\left[\frac{\phi^{2}}{2}\right]+\cdots\Bigr{\rangle}\,.$
Again, we may consider the above relation off-criticality and consider the $m^{2}$-expansion. The leading term in this expansion is determined by the OPE coefficients associated with the identity, which must satisfy
$c_{\phi\bar{\phi}1}\left(t,x\right)$ $=$ $-\frac{1}{D}\partial_{t}c_{\phi\phi 1}\left(t,x\right)\,.$
Moreover, since the non-analytic dependence on the mass is captured by the VEV of composite operators, one also obtains
$c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)$ $=$ $-\frac{1}{D}\partial_{t}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)\,.$
It is straightforward to check that the coefficients calculated in sections V.1 and V.2 satisfy the above non-perturbative relations.
VI Universal quantities from the OPE
In the previous section we obtained certain OPE coefficients. Now we shall see that they entail universal information after implementing suitable normalization conditions. The arguments used in this section follow the line of reasoning detailed in ref. Amit (1984), to which we refer for a review of universality in the context of the perturbative RG and its application to the statistical field theory approach for critical phenomena.
VI.1 Universality in the static limit
In critical systems the most calculated universal quantities are, by far, the critical exponents, which turn out to display universality in a rather direct way. There are, however, further quantities that display universality. In particular, the so-called universal amplitude ratios and the universal form of the equation of state have been studied Brezin et al. (1972, 1973, 1974); Bervillier (1976). In order to define such quantities, the implementation of some convention is required. For instance, in the case of the equation of state two normalizations must be imposed Amit (1984). In the case of of the universal equation of state, initially one calculates a certain expression, which must be consistent with the associated Callan-Symanzik (CS) equation. The analysis of the CS equation allows one to pinpoint where the non-universal contributions come from and it turns out that a scaling variable [^3] It can be a field or a coupling, such as the mass in the case of the universal equation of state. is typically multiplied by a factor that depends on the integrated trajectory along the RG flow rather than only on the fixed point itself. The idea is then to define one or more new variables by imposing certain conventions in such a way that this non-universal factor is “eaten up” by the new variables and one is left with a universal result.
In order to make contact with the OPE coefficients, let us start by considering the static case. In many cases of interest, the static limit is described not just by a scale-invariant theory but by a CFT. In the CFT framework, once a normalization is fixed for the operators, typically $\langle O\left(x\right)O\left(0\right)\rangle=1/x^{2\Delta_{O}}$, the OPE coefficients take the form $c_{ijk}/x^{\Delta_{i}+\Delta_{j}-\Delta_{k}}$ where the number $c_{ijk}$ is as universal as the scaling dimensions $\Delta_{i}$ since both are determined solely on the basis of symmetry and field content. Let us provide a brief argument from the field theory perspective via the following example: we consider the scaling field $\phi_{1}\equiv{\cal N}_{1}\phi$ and $\phi_{2}\equiv{\cal N}_{2}\left[\phi^{2}/2\right]$. [^4] The elementary field $\phi$ is a scaling operator. For a composite operator, one may have to take a suitable combination of composite operators. The normalizations ${\cal N}_{i}$ are determined by imposing $\langle\phi_{i}\left(x\right)\phi_{j}\left(0\right)\rangle=\delta_{ij}/x^{2\Delta_{i}}$. Next, we consider the ratio
$\frac{\langle\phi\left(p_{1}\right)\phi\left(p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi\left(p_{1}-p_{2}\right)\phi\left(-p_{1}+p_{2}\right)\rangle\langle\left[\frac{\phi^{2}}{2}\right]\left(p_{1}+p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$
It is straightforward to check that this ratio can be rewritten as follows
$\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$
The two-point function in the denominator are just the Fourier transform of our normalization condition $\langle\phi_{1}\left(p\right)\phi_{1}\left(-p\right)\rangle=c_{1}p^{\alpha_{1}}$ and $\langle\phi_{2}\left(p\right)\phi_{2}\left(-p\right)\rangle=c_{2}p^{\alpha_{2}}$ calculated for some given momenta. The striking feature of this ratio is that the associated CS equation does not depend on the wave functions renormalizations of $\phi$ and of $\left[\phi^{2}\right]$ implying that the following equation holds at zero mass:
$\left(\mu\partial_{\mu}+\beta\partial_{g}\right)\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}$ $=$ $0\,.$
By using the arguments detailed in Amit (1984), one can see that the above ratio is universal. Thus, by applying the OPE to the three-point function above it is straightforward to conclude that the coefficient $c_{112}$ is universal too.
VI.2 Universality in the dynamical case
VI.2.1 General discussion
Let us consider a two-point function $C\left(t,x\right)$, such as $\langle\phi\phi\rangle$ or $\langle\phi\bar{\phi}\rangle$. At a fixed point scale invariance implies that
$C\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta}}F\left(c_{1}\frac{t}{x^{z}}\right)\,,$
where $c_{1}$ is a non-universal constant. If $F\left(0\right)=c_{2}$ one can rewrite
$C\left(t,x\right)$ $=$ $\frac{c_{2}}{x^{\Delta}}f\left(c_{1}\frac{t}{x^{z}}\right)\,,$ (18)
where $f\left(0\right)=1$ and $c_{2}$ is a non-universal amplitude. The form of the function $f$ is universal and is fully specified once a further condition is imposed. This is needed since one could equally well consider a constant $\tilde{c}_{1}=\alpha c_{1}$, with $\alpha$ being a numerical factor. Therefore, we introduce the scaling variable $y={\cal N}_{y}t/x^{z}$ and write
$C\left(t,x\right)$ $=$ $\frac{c_{2}}{x^{\Delta}}f\left(c_{1}{\cal N}_{y}^{-1}y\right)\,\equiv\,\frac{c_{2}}{x^{\Delta}}f_{y}\left(y\right)\,,$
where ${\cal N}_{y}$ is determined by imposing a suitable condition, such as
$f_{y}\left(y=1\right)$ $=$ ${\rm const}.$
It is important to note that the scaling variable $y$ can be used in all two-point functions since they depend on the combination $c_{1}t/x^{z}$, where the non-universal constant $c_{1}$ is related to the RG running of the coupling $D$.
Let us now move to consider more specifically the OPE coefficients. At the fixed point, scale invariance implies that a OPE coefficient takes the form
$C_{ijk}\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta_{i}+\Delta_{j}-\Delta_{k}}}F_{ijk}\left(c_{1}\frac{t}{x^{z}}\right)\,.$ (19)
In the static limit, we recover the equilibrium case and the discussion of section VI.1 applies. Thus, one can still normalize the operators as discussed in section VI.1. For example, we define the scaling field $\phi_{1}\equiv{\cal N}_{1}\phi$, where ${\cal N}_{1}$ is determined by
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,.$
This normalization choice allows one to get rid of non-universal overall factors like $c_{2}$ in (18). However, even after this normalization, an OPE coeffient does not yet display universality away from the static limit due to the presence of $c_{1}$ in (19). Nevertheless, it is now straightforward to obtain a universal form of the OPE coefficient by moving to the normalized scaling variable $y$. Once expressed in this variable, the OPE coefficient of a normalized scaling operator will take the form
$C_{ijk}\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta_{i}+\Delta_{j}-\Delta_{k}}}c_{ijk}\left(y\right)\,,$
where $c_{ijk}\left(y\right)$ is a universal scaling function.
VI.2.2 Application to our results
We must first proceed to define suitably normalized scaling fields. We begin by considering $\phi_{1}\equiv{\cal N}_{1}\phi$ and imposing
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,,$
with $\Delta_{\phi}=\frac{d-2}{2}$. Within our approximations, this implies
${\cal N}_{1}$ $=$ $\left(\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\right)^{-1/2}\,.$
We then have
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}\,.$
The above two-point function is clearly non-universal as it depends on the non-universal coupling $D$. We introduce the scaling variable $y\equiv{\cal N}_{y}t/x^{z}$, where ${\cal N}_{y}$ is determined by
$x^{2\Delta_{\phi}}\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $1-\frac{\Gamma\left(\frac{d}{2}-1,1\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,.$
This choice is made for pure convenience and others are possible. We obtain ${\cal N}_{y}=4D$ and rewrite
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}g\left(y\right)\mbox{ with }g\left(y\right)=1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,,$
where the function $g\left(y\right)$ is universal.
Next, we wish to normalize the scaling field $\bar{\phi}_{1}\equiv\bar{{\cal N}}_{1}\phi$ by imposing a suitable normalization condition. Let us note that we cannot use the two-point $\langle\bar{\phi}_{1}\bar{\phi}_{1}\rangle$ since it vanishes. Therefore, we consider the response function $\langle\phi_{1}\left(t,x\right)\bar{\phi}_{1}\left(0,0\right)\rangle$ for $t>0$, so that it does not vanish due to the presence of $\theta\left(t\right)$. In particular we impose
$\langle\phi_{1}\left(t>0,x\right)\bar{\phi}_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $\frac{1}{x^{d}}\,,$
which implies $\bar{{\cal N}}_{1}=\frac{\pi^{d/2}e}{{\cal N}_{1}}$.
We are now ready to apply these normalizations to our results. Let us consider the OPE of the product $\phi_{1}\times\phi_{1}$:
$\phi_{1}\times\phi_{1}$ $=$ ${\cal N}_{1}^{2}\phi\times\phi\,=\,{\cal N}_{1}^{2}c_{\phi\phi 1}+{\cal N}_{1}^{2}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left[\frac{\phi^{2}}{2}\right]+\cdots$
$=$ ${\cal N}_{1}^{2}c_{\phi\phi 1}+{\cal N}_{1}^{2}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}{\cal N}_{2}^{-1}\phi_{2}+\cdots$
$=$ $c_{\phi_{1}\phi_{1}1}+c_{\phi_{1}\phi_{1}\phi_{2}}\phi_{2}+\cdots\,.$
Thus, we obtain
$c_{\phi_{1}\phi_{1}1}\left(t,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}}}\left\{1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\right\}+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{\phi_{2}}}}\sqrt{2}\left(1-\frac{\epsilon}{6}+\frac{\epsilon}{6}\Gamma\left(0,\frac{1}{y}\right)\right)+O\left(\epsilon^{2}\right)\,.$
In the present approximation $c_{\phi_{1}\phi_{1}1}\left(t,x\right)$ is fixed by the Gaussian two-point function together with the chosen normalization conventions. The coefficient $c_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ is more interesting: in the static limit, i.e., $y\rightarrow 0$, we recover the $\epsilon$-correction known from CFT techniques Gaiotto et al. (2014); Gopakumar et al. (2017a, b); Dey et al. (2017); Gopakumar and Sinha (2018). In the dynamical case, this universal coefficient is promoted to be a function of the scaling variable $y$, which begins to be non-trivial from $O\left(\epsilon\right)$.
In full analogy, we obtain the following results for the product $\phi_{1}\times\bar{\phi}_{1}$:
$c_{\phi_{1}\bar{\phi}_{1}1}\left(t,x\right)$ $=$ $\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}}}\frac{4e^{-1/y}y^{-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\bar{\phi}_{1}\phi_{2}}\left(t,x\right)$ $=$ $\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{\phi_{2}}}}\left(-\epsilon\frac{e^{1-\frac{1}{y}}y^{-1}}{3\sqrt{2}}\right)+O\left(\epsilon^{2}\right)\,.$
VII Extension to $O\left(N\right)$ models
In this section we generalize the previous results to the critical dynamics associated with $O\left(N\right)$ models. We consider the action
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}^{m}\left(t,x\right)\Biggr{\{}\partial_{t}\phi_{0}^{m}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}^{n}\left(t,x\right)\phi_{0}^{n}\left(t,x\right)\right)\phi_{0}^{m}\left(t,x\right)$
$-D_{0}\bar{\phi}_{0}^{m}\left(t,x\right)\bar{\phi}_{0}^{m}\left(t,x\right)\Biggr{\}}\,.$
In $d=4-\epsilon$ at one-loop, the fixed point value of the interaction coupling is given by $\epsilon\frac{3}{\left(N+8\right)}\left(4\pi\right)^{2}$ while the mass anomalous dimension is modified by a factor $\frac{N+2}{3}$ with respect to the $N=1$ case.
VII.1 OPE at the Gaussian level
It is instructive to inspect the OPE of two fields at the Gaussian level. Let us define the composite operators as normal ordered operators, in full analogy with section IV.2. One obtains
$\phi^{a}\times\phi^{b}$ $=$ $G_{ab}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]+\cdots\,.$ (20)
For later purposes it is convenient to consider the following combinations of operators of dimension $d-2$:
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}$$\equiv$ $\left[\frac{\phi^{a}\phi^{b}}{2}-\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]$
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$$\equiv$ $\left[\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]\,,$
where the Einstein summation convention is used. By using the above combination we can rewrite equation (20) as
$\phi^{a}\times\phi^{b}$ $=$ $G_{ab}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}+\cdots\,.$
At this stage the introduction of the “transverse” and “longitudinal” composite operators may seem arbitrary. However, as we shall see in the next section, they have different scaling dimensions and also their role in the OPE is different. This difference also appears in the OPE of $\phi^{a}\times\bar{\phi}^{b}$.
VII.2 Leading correction to the OPE in the mass expansion
In this section we compute certain OPE coefficients by considering the mass expansion of the two-point function. We follow the approach of section V, i.e., we shall consider the two-point function close to criticality and compare its mass expansion with the OPE. By doing so, we are able to compute the OPE coefficient associated with operators that have a non-zero VEV.
It is straightforward to check that $\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}$ has a vanishing VEV. Therefore, we expect that it is not possible to calculate the associated OPE coefficient by looking at the mass expansion of the two-point function. Thus, in the following we will focus mostly on $\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$. Since $\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$ is proportional to $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$, we shall work with the latter from now on.
VII.2.1 The composite operator $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$
In this section we focus on the VEV and on the normalization of $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$ and compare the result against the $N=1$ case of section III.1.
Before considering the VEV of $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$, let us mention that it is straightforward to compute the anomalous dimension of $\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$ and $\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}$, which reads $\gamma_{S}=\frac{N+2}{3}\frac{g}{\left(4\pi\right)^{2}}+O\left(\epsilon^{2}\right)$ and $\gamma_{T}=\frac{2}{3}\frac{g}{\left(4\pi\right)^{2}}+O\left(\epsilon^{2}\right)$, respectively. At the fixed point, the associated scaling dimensions then read $\Delta_{S}\approx 2-\frac{6\epsilon}{N+8}$ and $\Delta_{T}\approx 2-\frac{N+6}{N+8}\epsilon$, respectively. (See Dey et al. (2017) for a calculation based on the so-called analytic bootstrap.)
At order $O\left(g^{0}\right)$, the VEV of $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$ is equivalent to the sum of $N$ VEVs of the single field $\left[\frac{\phi^{2}}{2}\right]$ of the $N=1$ case. At order $O\left(g\right)$, a straightforward calculation shows that all the diagrams and counterterms are the same as the $N=1$ case but are multiplied by a factor $N\frac{N+2}{3}$.
The amplitude $\langle\left[\frac{\phi^{m}\phi^{m}}{2}\right]\left[\frac{\phi^{n}\phi^{n}}{2}\right]\rangle$ receives fully analogous modifications: the order $O\left(g^{0}\right)$ is equal to the $N$-times the $N=1$ case while the order $O\left(g\right)$ is equal to the $N\frac{N+2}{3}$-times the $N=1$ case.
VII.2.2 OPE coefficients $c_{\phi\phi\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$ and $c_{\phi\bar{\phi}\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$
To evaluate the coefficients $c_{\phi\phi\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$ and $c_{\phi\bar{\phi}\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$ from the mass expansion of the propagator we consider the two-point functions $\langle\phi^{a=1}\phi^{a=1}\rangle$ and $\langle\phi^{a=1}\bar{\phi}^{a=1}\rangle$, where we fixed the indices to $a=b=1$ without loss of generality. The two-point function and response function are given by the very same expression that was considered in section V and are given in Appendix A. The only difference comes from the fact that the mass correction is given by the substitution (16) with the order $O\left(g\right)$ term being multiplied by $\frac{N+2}{3}$. One then finds
$c_{\phi\phi\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$$\approx$ $\frac{1}{N}\Biggr{\{}2+g\frac{N+2}{3}\left(\frac{\log\left(\pi x^{2}\mu^{2}\right)+\gamma}{16\pi^{2}}+\frac{1}{\left(4\pi\right)^{2}}\Gamma\left(0,\frac{x^{2}}{4Dt}\right)\right)\Biggr{\}}\,,$ (21)
where we neglected corrections due to the mass and of higher order in the interaction coupling.
In full analogy, we can evaluate the coefficient associated with the response function, it gives
$c_{\phi\bar{\phi}\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$$\approx$ $-g\frac{N+2}{3N}\frac{\theta\left(t\right)}{\left(4\pi\right)^{d/2}}\frac{1}{\left(Dt\right)}e^{-\frac{x^{2}}{4Dt}}\,.$ (22)
Clearly, both expressions (21) and (22) go back to those already found for $N=1$.
In order to obtain a universal form of these results we must suitably rescale the operators and thereby the coefficients. We define normalization conditions exactly as in section VI.2.2. By doing so, one obtains
$c_{\phi_{1}\phi_{1}S}\left(y,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{S}}}\sqrt{\frac{2}{N}}\left(1-\frac{N+2}{2\left(N+8\right)}\epsilon+\frac{N+2}{2\left(N+8\right)}\Gamma\left(0,\frac{1}{y}\right)\epsilon\right)+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\bar{\phi}_{1}S}\left(y,x\right)$ $=$ $-\frac{1}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{S}}}\frac{\left(N+2\right)}{\sqrt{2N}\left(N+8\right)}\frac{e^{1-\frac{1}{y}}\theta\left(y\right)}{y}\epsilon+O\left(\epsilon^{2}\right)\,.$
The static limit of $c_{\phi_{1}\phi_{1}S}$ matches with the result found in Dey et al. (2017) via the analytic conformal bootstrap.
VIII Large N limit
We consider the dynamical $O\left(N\right)$ model in dimension $42<d<$ 4.
VIII.1 Action in the large N limit
In this section we rewrite the action of the model A in a way that allows one to take the large $N$ limit straightforwardly. In order to do so, we translate the SDE from Ito to the Stratonovich interpretation and consider the associated path integral. The reason for this is that certain formal manipulations are based on the standard rules of calculus, which are not always valid for the path integral. This is particularly relevant when there are “ordering ambiguities”, i.e., terms such as $\hat{p}\hat{q}$ in quantum mechanics. As we have seen in section IV.1, the Fokker-Planck Hamiltonian contains such terms. Ordering ambiguities are in one-to-one correspondence with the choice of path integral discretization Berezin (1980) and we shall stick with the midpoint prescription, which is associated with Weyl ordering. As reviewed in Appendix B, by using this discretization an integral over a time derivative satisfies the usual Stokes theorem.
To study the large $N$ limit we follow the approach explained in Moshe and Zinn-Justin (2003). The partition function takes the form
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$
×$\times\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}\,,$
where the last term arises from the non-trivial functional determinant present when using the midpoint discretization (see, e.g., Täuber (2014)). It reads
$\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}$ $=$ $\exp\Biggr{[}\frac{1}{2}D\int_{t,xy}\left(N\left(-\partial_{x}^{2}+r\right)+\frac{u_{0}}{3!}\left(\phi^{2}+\frac{2}{N}\phi^{2}\right)\right)\delta\left(x-y\right)\Biggr{]}\,.$
We rewrite the partition function by inserting a unit factor via a delta function:
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$
$\int{\cal D}\lambda{\cal D}\rho\,e^{i\lambda\left(\rho-\phi^{2}\right)}\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}$
$=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\rho\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$
$\int{\cal D}\lambda{\cal D}\rho\,e^{i\lambda\left(\rho-\phi^{2}\right)}\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\left(\rho\right)\Biggr{]}\,.$
Next, we split the fields: $\phi^{a}=\left(\sigma,\pi^{i}\right)$ and $\bar{\phi}^{a}=\left(\bar{\sigma},\bar{\pi}^{i}\right)$ and integrate over $\pi$ and $\bar{\pi}$. One obtains the following action
$S$ $=$ $\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\rho\right)\sigma\right)$
$-i\int_{tx}\lambda\left(\rho-\sigma^{2}\right)-\int_{tx}D_{0}\bar{\sigma}\bar{\sigma}\Biggr{]}-\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\left(\rho\right)$
$+\frac{N-1}{2}\mbox{Tr}\log\left(\frac{1}{2D_{0}}\left(\partial_{t}+D_{0}\left(\cdots\right)\right)\left(-\partial_{t}+D_{0}\left(\cdots\right)\right)+2i\lambda\right)\,.$
To make the large $N$ limit manifest we make some rescalings: $\left(\sigma,\bar{\sigma}\right)\rightarrow\sqrt{N}\left(\sigma,\bar{\sigma}\right)$, $\lambda\rightarrow\frac{\lambda}{2}$, and $\rho\rightarrow N\rho$ so that the action now reads
$S$ $=$ $N\Biggr{\{}\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma\right)$
$-i\int_{tx}\frac{\lambda}{2}\left(\rho-\sigma^{2}\right)-\int_{tx}D_{0}\bar{\sigma}\bar{\sigma}\Biggr{]}$
$-\frac{1}{2}D\int_{t,xy}\left(\left(-\partial_{x}^{2}+r\right)+\frac{u_{0}}{3!}\rho\right)\delta\left(x-y\right)$
$+\frac{1}{2}\mbox{Tr}\log\left(\frac{1}{2D_{0}}\left(\partial_{t}+D_{0}\left(\cdots\right)\right)\left(-\partial_{t}+D_{0}\left(\cdots\right)\right)+i\lambda\right)\Biggr{\}}\,,$
where we also used $N-1\approx N$ and we neglected subleading terms in the $1/N$ expansion of the functional determinant.
The functional integral is then evaluated by employing the saddle-point approximation. We look for constant solutions to the equation of motion, which we write below:
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma-2D_{0}\bar{\sigma}$
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\bar{\sigma}+i\lambda\sigma$
00 0 $=$ $-i\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{\omega q}\frac{i}{\frac{\omega^{2}+D_{0}^{2}\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)^{2}}{2D_{0}}+i\lambda}$
00 0 $=$ $\bar{\sigma}D_{0}\frac{u_{0}}{3!}\sigma-i\frac{\lambda}{2}-\frac{1}{2}D_{0}\left(\int_{q}\right)\frac{u_{0}}{3!}$
$+\frac{1}{2}\int_{\omega q}\frac{D_{0}\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)}{\left(\frac{\omega^{2}+D_{0}^{2}\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)^{2}+2D_{0}i\lambda}{2D_{0}}\right)}\frac{u_{0}}{3!}\,.$
In the present approximation, the constant solutions represent the expectation values of the fluctuating fields themselves. Thus, the solution for $\bar{\sigma}$ approximate $\langle\bar{\sigma}\rangle$. Since we know that $\langle\bar{\sigma}\rangle=0$ we look for solutions such that $\bar{\sigma}=0$. Furthermore, since these constant solutions represent expectation values of the fields, which are time-independent due to time translation invariance, we expect that they satisfy the same equation of the static case. It is possible to check that for $\lambda=0$ one recovers the static equations. In particular, the second and the fourth equations vanish identically and one is left with
00 0 $=$ $\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma$
00 0 $=$ $-\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)}\,.$
It is convenient to define $m^{2}\equiv r_{0}+\frac{u_{0}}{3!}\rho$ so that
00 0 $=$ $m^{2}\sigma$
00 0 $=$ $\frac{\sigma^{2}}{2}-\frac{3}{u}\left(m^{2}-r_{0}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+m^{2}\right)}\,.$
In the symmetric phase $\sigma=0$ and $m^{2}\neq 0$ while for the broken phase $\sigma\neq 0$ and $m^{2}=0$. By employing dim-reg, the symmetric phase value of $m^{2}$ is determined by
00 0 $=$ $\frac{3!}{u_{0}}\left(r_{0}-m^{2}\right)+\frac{\Gamma\left(1-d/2\right)}{\left(4\pi\right)^{d/2}}\left(m^{2}\right)^{\frac{d-2}{2}}\,.$
In this case the critical value of the mass is $r_{0,c}=0$ and $m^{2}=0$. In the symmetric phase and very close to criticality we can approximate
$m^{2}$$\approx$ $\left(-\frac{\left(4\pi\right)^{d/2}}{\Gamma\left(1-d/2\right)}\frac{3!}{u_{0}}r_{0}\right)^{\frac{2}{d-2}}\,\equiv\,m_{{\rm o.s.}}^{2}\,.$
Note that this allows one to compute straightforwardly the expectation value $\langle\phi^{2}\rangle=\langle N\rho\rangle\approx N\frac{3!}{u_{0}}\left(m_{{\rm o.s.}}^{2}-r_{0}\right)$.
VIII.2 OPE coeffients
To compute the OPE coefficient, we consider the mass-expansion of the two-point function. In order to identify the OPE coefficients, we assume that in this expansion the non-analytic dependence on $r_{0}$ is due to the expectation value of composite operators. [^5] The non-analytic dependence on $u_{0}$ is instead expected from the form that the action can take in the static limit, see Rose et al. (2022) and references therein.
In the large $N$ limit one obtains
$\langle\sigma\sigma\rangle$$\approx$ $\frac{1}{N}\frac{1}{\left(4\pi\right)^{d/2}}\Biggr{\{}2^{d-2}\Gamma\left(\frac{d}{2}-1\right)\frac{1}{x^{d-2}}+\Gamma\left(1-\frac{d}{2}\right)m_{{\rm o.s.}}^{d-2}-2^{d-4}\Gamma\left(\frac{d}{2}-2\right)x^{4-d}m_{{\rm o.s.}}^{2}$
$-\frac{2^{d-2}}{x^{d-2}}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)+\frac{2^{d-4}}{x^{d-2}}m_{{\rm o.s.}}^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4Dt}\right)\Biggr{\}}\,.$
The part non-analytic in $r_{0}$ is due to $m_{{\rm o.s.}}^{2}$, which is related to $\frac{u_{0}}{3!N}\langle\phi^{2}\rangle$. Thus, we can read off the OPE coefficient and normalize the operator in the same way as we did in section VI. The normalization of $\phi_{2}$ has already been computed in the static theory in Rose et al. (2022), to which we refer for the detailed calculation of the normalization. To the leading order in $1/N$, one finally obtains
$c_{\phi_{1}\phi_{1}\phi_{2}}$ $=$ $\frac{1}{x^{2\Delta_{1}-\Delta_{2}}}\sqrt{\frac{2}{N}}\sqrt{\frac{\Gamma\left(d-2\right)\sin\left(\pi\frac{d}{2}\right)}{\left(\frac{d}{2}-2\right)\pi}}\frac{1}{\Gamma\left(\frac{d}{2}-1\right)}\left[1-\frac{\Gamma\left(\frac{d}{2}-2,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-2\right)}\right]\,.$
In full analogy, one can determine the OPE coefficient appearing in the response function. After the normalization of the operators it takes the form
$c_{\phi_{1}\bar{\phi}_{1}\phi_{2}}$ $=$ $-\frac{2}{\sqrt{N}\sqrt{\pi}}\frac{\theta\left(y\right)}{x^{d-2}}e^{1-\frac{1}{y}}y^{1-\frac{d}{2}}\sqrt{\frac{\sin\left(\frac{\pi d}{2}\right)\Gamma\left(d-2\right)}{\left(d-4\right)\Gamma\left(\frac{d}{2}-2\right)^{2}}}\,.$
to the leading order in $1/N$.
IX Summary and conclusions
In this work we considered the OPE in the context of critical dynamics and computed some OPE coefficients for the so-called model A. Moreover, we showed how it is possible to compute certain universal quantities that are related to the OPE coefficients by some suitable normalizations. The role of the fluctuation-dissipation theorem for the calculated coefficients has also been discussed.
More precisely, in section IV we derived the Gaussian form of the OPE. We found it instructive to derive it by introducing a normal ordering prescription stemming from interpreting the Fokker-Planck equation as a functional “Schroedinger” equation. The corrections to the Gaussian OPE have been calculated in section V by considering a mass expansion of the correlation function of interest. By defining suitable normalizations we have derived a universal form of the OPE coefficients in section VI. The $O\left(N\right)$ dynamical model has been considered in sections VII and VIII where the perturbation theory and the large $N$ limit have been considered.
It is possible to extend the present study in several directions. On the one hand, it is possible to push our calculations to higher orders and to consider further OPE coefficients. On the other hand, the present paper considers the critical dynamics associated with systems that have already relaxed to equilibrium. Therefore, it would be interesting to generalize our study also to the case of non-thermal fixed points Berges (2015) or to consider the effect of the initial state in more detail Täuber (2014).
Acknowledgments
C. P. thanks prof. Andrea Gambassi for useful discussions.
Appendix A Conventions
We employ the following short-hand notations. In real space
[…]
while in momentum space
$\int_{\omega}\equiv\int\frac{d\omega}{2\pi}$ and $\int_{q}\equiv\int\frac{d^{d}q}{\left(2\pi\right)^{d}}\,.$
The Fourier and inverse Fourier transforms are defined as follows:
$f\left(t,x\right)$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\tilde{f}\left(\omega,q\right)$
$\tilde{f}\left(\omega,q\right)$ $=$ $\int_{tx}e^{+i\omega t-iqx}f\left(t,x\right)\,.$
The two-point function in the free-field theory approximation are given by the following expressions:
$\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\frac{2D}{\omega^{2}+D^{2}\left(q^{2}+m^{2}\right)^{2}}$ (23)
$=$ $\frac{1}{\left(4\pi\right)^{d/2}}2^{d/2}\frac{1}{x^{d-2}}\left(m^{2}x^{2}\right)^{\frac{d-2}{4}}K_{\frac{d-2}{2}}\left(\sqrt{m^{2}x^{2}}\right)$
[…]
$=$ $\frac{1}{\left(4\pi\right)^{d/2}}\frac{1}{x^{d-2}}\Biggr{\{}2^{d/2}\left(m^{2}x^{2}\right)^{\frac{d-2}{4}}K_{\frac{d-2}{2}}\left(\sqrt{m^{2}x^{2}}\right)$
$-2^{d-2}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)+2^{d-4}m^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4Dt}\right)\Biggr{\}}+o\left(m^{2}\right),$
and
$\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$ $=$ $\int_{\omega}\int_{q}e^{-i\omega t}e^{iqx}\frac{1}{-i\omega+D\left(q^{2}+m^{2}\right)}$ (24)
$=$ $\theta\left(t\right)\frac{1}{\left(4\pi\right)^{d/2}}\frac{1}{\left(Dt\right)^{d/2}}e^{-m^{2}Dt-\frac{x^{2}}{4Dt}}\,.$
Appendix B Fokker-Planck Hamiltonian, ordering, and mid-point prescription
For the purposes of this appendix, it suffices to consider a generic one-dimensional FP equation:
$\partial_{t}p\left(x,t\right)$ $=$ $-\partial_{x}\left(A\left(x\right)p\left(x,t\right)\right)+D\partial_{x}^{2}p\left(x,t\right)\,.$
By introducing $\hat{p}\equiv-i\partial_{x}$ we can rewrite
$\partial_{t}p\left(x,t\right)$ $=$ $\left[-i\hat{p}A\left(\hat{x}\right)-D\hat{p}^{2}\right]p\left(x,t\right)$
$=$ $-\left[i\hat{p}A\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$
As it is written, the ordering is of the $\hat{p}\hat{x}$-type and if one proceeds to construct the associated path integral one obtains the “Ito action”, which is assiociated with the pre-point discretization. The same FP equation can be rewritten by introducing
$\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}$$\equiv$ $\frac{\hat{p}A\left(\hat{x}\right)+A\left(\hat{x}\right)\hat{p}}{2}\,.$
One finds
$\partial_{t}p\left(x,t\right)$ $=$ $-\left[i\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}+\frac{1}{2}A^{\prime}\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$
If one constructs the path integral associated with the above equation the associated action is composed by a part which is fully similar to the “Ito-action” (but whose associated discretization is dictated by the midpoint rule) and by a new term, which corresponds to the second term in the brackets. Such term precisely gives rise to the well-known determinant. Thus, this clarifies the presence of the determinant when using the midpoint discretization.
Let us now discuss the role of the midpoint discretization when dealing with surface terms, see Chaichian and Demichev (2001). In particular, one wishes to use the standard relation
[…]
If one takes the discretized point to correspond to $\left(1-\lambda\right)x_{j}+\lambda x_{j+1}$ with $\lambda\in\left[0,1\right]$ one recovers the prepoint case for $\lambda=0$ and the midpoint case for $\lambda=1/2$. By denoting $\delta\equiv x_{j+1}-x_{j}$, the above surface term can be expressed as follows
$f\left(x\left(t_{f}\right)\right)-f\left(x\left(t_{0}\right)\right)\,=\,\sum_{j}\left[f\left(x_{j+1}\right)-f\left(x_{j}\right)\right]$ $=$
$\sum_{j}f^{\prime}\left(x_{j}+\lambda\delta\right)\delta+\sum_{j}\frac{1}{2}\left(1-2\lambda\right)f^{\prime\prime}\left(x_{j}\right)\delta^{2}+O\left(\delta^{3}\right)\,.$
For the midpoint choice, i.e., $\lambda=1/2$ the second term does not appear. However, such terms are generally relevant in the path integral. Indeed we know that the path integral is determined by the paths which are continuous but nowhere differentiable implying that $\delta^{2}\sim\Delta t$ and so that the second term gives rise to a finite term in the action in the continuum limit. To avoid dealing with such terms we adopt the midpoint discretization.